one leg of a right triangle is 6cm shorter than the other leg. the length of the hypotenuse is 30cm. find the length of the longer leg.
Answer by stanbon(57282)

one leg of a right triangle is 6cm shorter than the other leg. the length of the hypotenuse is 30cm. find the length of the longer leg.
---------------------
one leg: x
other leg: x-6
hyp = 30
----------------------------------
Pythagoras:
x^2 + (x-6)^2 = 30^2
x^2 + x^2-12x+36 = 900
2x^2 -12x - 864 = 0
x^2 - 6x - 432 = 0
(x+18)(x-24) = 0
Positive solution:
x = 24 cm (one leg)
x-6 = 18 cm (other leg)
=========================
Cheers,
Stan H.